# SOFIA/HAWC+ Far-Infrared Polarimetric Large Area CMZ Exploration Survey. V. The Magnetic Field Strength and Morphology in the Sagittarius C Complex

### v1

@@ 4.2.1 Structure of Clouds in Sgr C
Figure 8: Velocity channel maps of the HCN J=1-0 line measured by the 3 mm Mopra survey (Jones et al., 2012). The survey has spatial and spectral resolutions of 36″ and 1.84 km s-1, respectively. The mean velocity of each channel is indicated in the top left corner. These channel maps cover the velocity range between $-92.6$ and $-34.3$ km s-1 in steps of $7.3$ km s-1. The grey-scale contours show the peak emission of HCN integrated through all observed velocities drawn at 0.75, 1.25, 1.75 K km s-1, with the brighter curves indicating higher intensity. The yellow schematics indicate the approximate placement and extent of the Sgr C H II region and the Sgr C NTF. The green star in the central panel denotes the location of the G359.44-0.10 EGO, which is addressed in detail in Section 5.3. The global low intensity in channels near $-50$ km s-1 is due to absorption in the foreground “3 kpc arm” (Lang et al., 2010). The 214 µm magnetic field pseudovectors in this figure are rendered using Nyquist sampling at the Mopra beam size.
@@
In the southern portion of the $-65$ km s-1 cloud, there is a bright bulk of HCN emission at $l=359.47$, $b=-0.17$ visible at all displayed velocities. In this region, we do not find any morphological relationship between the HCN emission and the 214 µm magnetic field orientations. However, many 850 µm magnetic field detections were reported in this cloud by Lu et al. (2024), and they do not generally agree with the 214 µm results of Paré et al. (2024). The HCN emission offers several possible explanations for this discrepancy. It could be that this is a particularly disturbed and turbulent cloud with multiple temperature components, or there could be multiple clouds superposed along the same line of sight. Both can cause the magnetic field measured from different wavelengths to exhibit different orientations. It is also likely that this cloud has a radial velocity gradient along the line of sight, given that this is where the rotating 100 pc ring reverses direction in projection. The observation of this particularly broad-linewidth cloud re-affirms our caveat for the DCF analysis in Section 3.2.4 regarding the incomplete accounting for multiple line-of-sight clouds having different velocities in the derivation of the magnetic field strengths.
The previously mentioned 214 µm magnetic field measurement cavity at $l=359.44$, $b=-0.13$ between the $-65$ km s-1 cloud and Cloud #4 (centered at $l=359.43$, $b=-0.17$; see Figure 4) is of interest to be studied. In Figure 8, the $-65$ km s-1 cloud and Cloud #4 appear to be portions of a continuous structure so they are likely physically related. We hypothesize that they started as one cloud and were later separated into the currently observed shape by an outside force that created the magnetic field measurement cavity at $l=359.44$, $b=-0.13$. We note a potentially relevant X-ray association with this cavity in Section 5.2.2.
@@ 4.2.2 Shock Tracer in the Twisted 100 pc Ring
As previously mentioned, Sgr C is located on the western vertex of a dust ring (Molinari et al., 2011), and is postulated to be a collision site of $X_{1}$-$X_{2}$ orbits due to the bar potential (Sormani et al., 2018, 2020), which likely induces shocks and star formation. Meanwhile, the expanding Sgr C H II region can also induce a compression front that collides with surrounding molecular clouds (Liszt & Spiker, 1995) and add to the effect of the large-scale shocks. We employ the many 214 µm magnetic field detections toward the molecular clouds in the Sgr C region to examine how such shocks might be manifested in the magnetic field orientations. Figure 9 shows the velocity channel maps of the SiO J=2-1 line, a known shock tracer also from the Mopra survey (Jones et al., 2012).
Figure 9: Velocity channel maps of SiO J=2-1 emission between $-110$ and $-49$ km s-1, measured by the Mopra survey (Jones et al., 2012). The 214 µm magnetic field pseudovectors are superimposed as white line segments, as in Figure 3. The yellow schematic indicates the approximate location and extent of the Sgr C H II region and the Sgr C NTF.
The SiO channel maps exhibit a similar velocity structure to that of HCN, indicating that the SiO emissions are from within the 100 pc ring. However, the SiO emission appears much less spatially extended than that of HCN and more concentrated around the Sgr C H II region. This local enhancement of SiO indicates that the portions of the 100 pc ring surrounding the Sgr C H II region are particularly subjected to shocks, likely due to a combination of 1) the impact of the dust lane inflowing along the Galactic bar upon the CMZ material in the 100 pc ring, and 2) the collision between ambient gas and the expanding shell surrounding the H II region (Riquelme et al., In-preparation). The collision likely increases the turbulence within the 100 pc ring, ultimately leading to a low polarization fraction, a higher turbulent-to-ordered magnetic field strength ratio, and relatively low inferred sky-plane magnetic field strength, as shown in Figure 5.
@@ 5.1 Associated NTFs & Radio Features
As has previously been widely discussed, NTFs are valuable probes of the large-scale magnetic field in the inter-cloud volume of the central few hundred parsecs of the Galaxy (Serabyn & Morris, 1994; Morris, 2006; Thomas et al., 2020). The following subsections examine the magnetic field behaviour around the NTFs and radio structures observed in the Sgr C region, keeping in mind that the FIREPLACE detections primarily probe the magnetic field within clouds.
@@ 5.1.1 The Sgr C NTF
Figure 10: The 214 µm magnetic field pseudovectors near the Sgr C NTF and G359.43+0.02 YSO superimposed on emission at 20 cm (Heywood et al., 2022), 8 µm (Stolovy et al., 2006), 24 µm (Carey et al., 2009), 250 µm (Molinari et al., 2010), and 850 µm (Pierce-Price et al., 2000). The yellow segments are the 214 µm magnetic field pseudovectors from Paré et al. (2024) and the orange segments are the 850 µm magnetic field pseudovectors from Lu et al. (2024).
This extended NTF appears luminous and compact below $l=-0.03$ and it splits into two parallel filaments and becomes fainter and more diffuse above this latitude. This NTF overlaps on the sky-plane with three distinct magnetic field intervals:
@@
Interval (ii) above is roughly the extent of Part A as defined in Roy (2003) and Lang et al. (2010). We note additionally that the Sgr C NTF has a faint end towards the south, detailed in Section 5.1.5.
In the radio continuum below $l=-0.02$, the NTF has the same width but Part A is more luminous than Part B (the part of the Sgr C NTF below $l=-0.08$ labelled in Figure 10; Roy 2003). The radio spectral index analysis of Yusef-Zadeh et al. (2022a) indicates that Part A has an average spectral index of $\sim-1.02$ and Part B has an average spectral index of $\sim-0.65$. In Part A, the magnetic field orientation near the NTF roughly follows its direction. As our magnetic field measurements sample the field from relatively cold clouds, the direction of the FIREPLACE magnetic field in the $-90$ km s-1 cloud is consistent with the orientation of the Sgr C NTF, as are the 850 µm measurements of Lu et al. (2024). Moreover, the DCF analysis from Figure 5 shows that there is a “tunnel” of low turbulent-to-ordered ratio and high $|{\bf B_{\rm POS}}|$ along the Sgr C NTF. However, such alignment ceases beyond $l>-0.02$ due to an absence of magnetic field detections outside the molecular cloud. Additionally, the NTF becomes much more diffuse after departing from the molecular cloud toward the north. Above $l=0$, the magnetic field orientation and the Sgr C NTF are no longer aligned, which could be attributed to strong far-IR emission from FIR-4 dominating the magnetic field detections in that region.
@@
At the northern portion of the Sgr C NTF where it splits into two filaments and begins to fade out toward the north, lies a short, linear radio feature oriented parallel to the Galactic plane along $b\approx 0.01$. It was originally found and denoted as FIR-4 by Odenwald & Fazio (1984) in the course of their far-IR survey of the GC. FIR-4 and its counterparts are displayed in Figure 10. The mid-IR counterpart of FIR-4 is the G359.43+0.02 YSO Cluster reported by Yusef-Zadeh et al. (2009) and prominently displayed in the central panel of Figure 10.
@@
Figure 11: The G359.33-0.15 NTF towards the southwest of the Sgr C H II region shown at 20 cm (Heywood et al., 2022; Yusef-Zadeh et al., 2022a), 8 µm (Stolovy et al., 2006), 24 µm (Carey et al., 2009), and 70 µm (Molinari et al., 2010), overlaid with 214 µm magnetic field pseudovectors from Paré et al. (2024). The orange segments in the lower right panel are the 850 µm magnetic field detections from Lu et al. (2024).
@@
Unlike most NTFs in the CMZ, which are perpendicular to the Galactic plane, the G359.33-0.15 filaments are much closer to being parallel to the Galactic plane. Another of its notable characteristics is that it is apparently interacting with an infrared dark cloud (IRDC) at its southwestern end (compare the top two panels in Figure 11 at near $l=359.32$, $b=-0.17$), and that interaction has led to a strong morphological bunching and brightening of the radio emission at the interface (see Fig. 3k of Yusef-Zadeh et al. (2022b) for additional radio detail). The interaction has also led to the production of a slightly extended X-ray source at or very near that location (Yusef-Zadeh et al., 2007, their designation: G359.32-0.16), raising the possibility that this location is where the relativistic electrons that illuminate the G359.33-0.15 filament are produced, by analogy with G359.1-0.2 (Uchida et al., 1996) or G0.18-0.04, and the Sickle (Serabyn & Morris, 1994). This interesting structure clearly warrants further investigation. Unfortunately, no magnetic field detections have been measured along the G359.33-0.15 NTF and the IRDC with which it is interacting due to a low far-IR luminosity (Paré et al., 2024).
@@
Figure 12: The FIREPLACE 214 µm magnetic field pseudovectors around Source C, superimposed on emission maps at 20 cm (Heywood et al., 2022), 8 µm (Stolovy et al., 2006), 24 µm (Carey et al., 2009), and 250 µm (Molinari et al., 2010). The yellow segments are the 214 µm magnetic field pseudovectors reported by Paré et al. (2024). (No 850 µm magnetic field detections from Lu et al. (2024) were reported near Source C.)
@@
Source C contains a short radio filament with a bright southern end that ends abruptly at $l=359.390$, $b=-0.077$ in the MeerKAT/L-Band image. To the northeast of that point, there are two adjacent and parallel filaments, the upper one of which appears to undergo a sudden northward deflection at $l=359.405$, $b=-0.06$, above which its brightness increases. The deflection point of this filament is located within a far-IR emission cavity centered at $l=359.405$, $b=-0.060$, where we find no significant magnetic field detection, but north of the cavity, the magnetic field orientation is approximately aligned with the filament. If the filament is a magnetic NTF, its “deflection” of the top filament could occur where it transitions going northward from the far-IR cavity into the local molecular cloud, thereby reflecting the differing field orientations in the two media.
@@
Figure 13: Top left: The full extent of the Sgr C NTF in radio continuum (Heywood et al., 2022). The orange contours are the FIREPLACE 214 µm emission at (3, 5.25, 7.5, 9.75, 12) Jy/pixel from Paré et al. (2024). Note the faint southern extension of the bright NTF bundle to the north. Top right: The faint southern “tail” of the Sgr C NTF, and source G359.467-0.17, displayed at wavelengths of 20 cm (Heywood et al., 2022), 8 µm (Stolovy et al., 2006), 24 µm (Carey et al., 2009), and 70 µm (Molinari et al., 2010), overlaid with the FIREPLACE 214 µm magnetic field pseudovectors from Paré et al. (2024). Bottom: The extended 20 cm radio emission towards the south of the Sgr C H II region, from Heywood et al. (2022), 8 µm (Stolovy et al., 2006), 250 µm (Molinari et al., 2010), and 850 µm (Pierce-Price et al., 2000) overlaid with the 214 µm magnetic field pseudovectors from Paré et al. (2024).
@@ 5.2 The Sgr C H II Region
To shed more light on the properties of the Sgr C H II region, we use the [C II] line analysis conducted by Riquelme et al. (In-preparation) using the upgraded German Receiver for Astronomy at Terahertz Frequencies (upGREAT) instrument aboard the SOFIA telescope. The [C II] line traces the warm ionized and partially ionized media exposed to UV photons from the HII region. Figure 14 displays the C II velocity channel image at $-41$ km s-1 (Riquelme et al., In-preparation) overlaid by the FIREPLACE 214 µm magnetic field pseudovectors for comparison with the morphology of the C II shell.
Figure 14: The CMZ in [C II] Survey image of the $-41$ km s-1 velocity channel observed using SOFIA/upGREAT (Riquelme et al., In-preparation). White segments are the 214 µm pseudovectors made from data reported by Paré et al. (2024). Pink stars indicate the positions of two Wolf-Rayet stars reported by Geballe et al. (2019). The cyan schematics mark the approximate location and extent of the Sgr C H II region and the Sgr C NTF.
In the immediate vicinity of the Sgr C H II region, the 850 µm magnetic field measurements reported by Lu et al. (2024) are largely consistent with the 214 µm ones shown here.
Figure 14 reveals what appears to be a double-cavity morphology associated with the Sgr C H II region, with the western cavity being larger in size. The [C II] shell surrounds and overlaps with the outer perimeter of the H II region appearing in radio emission (see Panel #9 in Figure 7), presumably identifiable as the ionization front. The 214 µm magnetic field orientation is mostly tangential to the cavities, except in the southeast direction because of the dense foreground head-tail cloud. The [C II] emission in this shell is found to deviate somewhat from circular symmetry, with the strongest emission in the northwest and near the G359.44-0.10 EGO. Together with the tangential magnetic field morphology, these observations resonate with the conclusion found in the simulations of Arthur et al. (2011), that the expansion of an H II region creates a tangential magnetic field orientation around its border. Moreover, the [C II] cavities are separated by a “bridge” of faint [C II] emission centered at $l=359.455$, $b=-0.090$. Only two significant 214 µm pseudovectors are present toward the “bridge,” and both are parallel to it.

### v2

@@ 4.2.1 Structure of Clouds in Sgr C
Figure 8: Velocity channel maps of the HCN J=1-0 line measured by the 3 mm Mopra survey (Jones et al., 2012). The survey has spatial and spectral resolutions of 36″ and 1.84 km s-1, respectively. The mean velocity of each channel is indicated in the top left corner. These channel maps cover the velocity range between $-92.6$ and $-34.3$ km s-1 in steps of $7.3$ km s-1. The grey-scale contours show the peak emission of HCN integrated through all observed velocities drawn at 0.75, 1.25, 1.75 K km s-1, with the brighter contours indicating higher intensity. The white line segments are the 214 µm FIREPLACE magnetic field pseudovectors from Paré et al. (2024) and the blue segments are the 850 µm magnetic field pseudovectors from Lu et al. (2024). The yellow schematics indicate the approximate placement and extent of the Sgr C H II region and the Sgr C NTF. The green star in the central panel denotes the location of the G359.44-0.10 EGO, which is addressed in detail in Section 5.3. The global low intensity in channels near $-50$ km s-1 is due to absorption in the foreground “3 kpc arm” (Lang et al., 2010). The 214 µm magnetic field pseudovectors in this figure are rendered using Nyquist sampling at the Mopra beam size.
@@
In the southern portion of the $-65$ km s-1 cloud, there is a bright bulk of HCN emission at $l=359.47$, $b=-0.17$ visible at all displayed velocities. In this region, we do not find any morphological relationship between the HCN emission and the 214 µm magnetic field orientations. However, many 850 µm magnetic field detections were reported in this cloud by Lu et al. (2024), and they do not generally agree with the 214 µm results of Paré et al. (2024). The HCN emission offers several possible explanations for this discrepancy. It could be that this is a particularly disturbed and turbulent cloud with multiple temperature components, or there could be multiple clouds superposed along the same line of sight, as we previously suggested in Section 3.2.4. Both can cause the magnetic field measured from different wavelengths to exhibit different orientations. It is also likely that this cloud has a radial velocity gradient along the line of sight, given that this is where the rotating 100 pc ring reverses direction in projection. The observation of this particularly broad-linewidth cloud re-affirms our caveat for the DCF analysis in Section 3.2.4 regarding the incomplete accounting for multiple line-of-sight clouds having different velocities in the derivation of the magnetic field strengths.
The previously mentioned 214 µm magnetic field measurement cavity at $l=359.44$, $b=-0.13$ between the $-65$ km s-1 cloud and Cloud #4 (centred at $l=359.43$, $b=-0.17$; see Figure 4) is of interest to be studied. In Figure 8, the $-65$ km s-1 cloud and Cloud #4 appear to be portions of a continuous structure so they are likely physically related. We hypothesize that they started as one cloud and were later separated into the currently observed shape by an outside force that created the magnetic field measurement cavity at $l=359.44$, $b=-0.13$. We note a potentially relevant X-ray association with this cavity in Section 5.2.2.
@@ 4.2.2 Shock Tracer in the Twisted 100 pc Ring
As mentioned in the introduction, Sgr C is located on the western vertex of a dust ring (Molinari et al., 2011), and is postulated to be a collision site of $X_{1}$-$X_{2}$ orbits due to the bar potential (Sormani et al., 2018, 2020), which likely induces shocks and star formation. Meanwhile, the expanding Sgr C H II region can also induce a compression front that collides with surrounding molecular clouds (Liszt & Spiker, 1995) and add to the effect of the large-scale shocks. We employ the many 214 µm magnetic field detections toward the molecular clouds in the Sgr C region to examine how such shocks might be manifested in the magnetic field orientations. Figure 9 shows the velocity channel maps of the SiO J=2-1 line, a known shock tracer also from the Mopra survey (Jones et al., 2012).
Figure 9: Velocity channel maps of SiO J=2-1 emission between $-110$ and $-49$ km s-1, measured by the Mopra survey (Jones et al., 2012). The 214 µm and 850 µm magnetic field pseudovectors are superimposed as white and blue line segments, as in Figure 8. The yellow schematic indicates the approximate location and extent of the Sgr C H II region and the Sgr C NTF.
The SiO channel maps exhibit a similar velocity structure to that of HCN, indicating that the SiO emissions are from within the 100 pc ring. However, the SiO emission appears much less spatially extended than that of HCN and more concentrated around the Sgr C H II region. This local enhancement of SiO indicates that the portions of the 100 pc ring surrounding the Sgr C H II region are particularly subjected to shocks, likely due to a combination of 1) the impact of the dust lane inflowing along the Galactic bar upon the CMZ material in the 100 pc ring, and 2) the collision between ambient gas and the expanding shell surrounding the H II region (Riquelme-Vásquez et al., 2025, submitted). The collision likely increases the turbulence within the 100 pc ring, ultimately leading to a low polarization fraction, a higher turbulent-to-ordered magnetic field strength ratio, and relatively low inferred sky-plane magnetic field strength, as shown in Figure 5.
@@ 5.1 Associated NTFs & Radio Features
NTFs are valuable probes of the large-scale magnetic field in the inter-cloud volume of the central few hundred parsecs of the Galaxy (Serabyn & Morris, 1994; Morris, 2006; Thomas et al., 2020). The following subsections examine the magnetic field behaviour around the NTFs and radio structures observed in the Sgr C region, keeping in mind that the FIREPLACE detections primarily probe the magnetic field within clouds.
@@ 5.1.1 The Sgr C NTF
Figure 10: The 214 µm magnetic field pseudovectors near the Sgr C NTF and G359.43+0.02 YSO superimposed on emission at 20 cm (Heywood et al., 2022), 8 µm (Stolovy et al., 2006), 24 µm (Carey et al., 2009), 250 µm (Molinari et al., 2010), and 850 µm (Pierce-Price et al., 2000). The yellow segments are the 214 µm magnetic field pseudovectors from Paré et al. (2024) and the orange segments are the 850 µm magnetic field pseudovectors from Lu et al. (2024). Important features in this region, such as FIR-4, the two parts of the Sgr C NTF, and the 90 km s-1 cloud, are labelled by orange arrows in different panels where the mentioned structures are visible.
This extended NTF appears luminous and compact below $l=-0.02$ and it splits into two parallel filaments and becomes fainter and more diffuse above this latitude. This NTF overlaps on the sky-plane with three distinct magnetic field intervals:
@@
Interval (ii) above is roughly the extent of Part A as defined in Roy (2003) and Lang et al. (2010) (also labelled in the leftmost panel of Figure 10). We note additionally that the Sgr C NTF has a faint end towards the south, detailed in Section 5.1.5.
In the radio continuum below $l=-0.02$ (indicated by the upper dashed line in the leftmost panel of Figure 10), the NTF has the same width but Part A is more luminous than Part B (the part of the Sgr C NTF below $l=-0.08$ labelled in Figure 10; Roy 2003). The radio spectral index analysis of Yusef-Zadeh et al. (2022a) indicates that Part A has an average spectral index of $\sim-1.02$ and Part B has an average spectral index of $\sim-0.65$. In Part A, the magnetic field orientation near the NTF roughly follows its direction. As our magnetic field measurements sample the field from relatively cold clouds, the direction of the FIREPLACE magnetic field in the $-90$ km s-1 cloud (labelled in the fourth panel in Figure 10) is consistent with the orientation of the Sgr C NTF, as are the 850 µm measurements of Lu et al. (2024). Moreover, the DCF analysis from Figure 5 shows that there is a “tunnel” of low turbulent-to-ordered ratio and high $|{\bf B_{\rm POS}}|$ along the Sgr C NTF. However, such alignment ceases beyond $l>-0.02$ due to an absence of magnetic field detections outside the molecular cloud. Additionally, the NTF becomes much more diffuse after departing from the molecular cloud toward the north. Above $l=0$, the magnetic field orientation and the Sgr C NTF are no longer aligned, which could be attributed to strong far-IR emission from FIR-4 dominating the magnetic field detections in that region.
@@
At the northern portion of the Sgr C NTF where it splits into two filaments and begins to fade out toward the north, lies a short, linear radio feature oriented parallel to the Galactic plane along $b\approx 0.01$, labelled in the leftmost panel of Figure 10). It was originally found and denoted as FIR-4 by Odenwald & Fazio (1984) in the course of their far-IR survey of the GC. FIR-4 and its counterparts are displayed in Figure 10. The mid-IR counterpart of FIR-4 is the G359.43+0.02 YSO Cluster reported by Yusef-Zadeh et al. (2009) and prominently displayed in the central panel of Figure 10.
@@
Figure 11: The G359.33-0.15 NTF towards the southwest of the Sgr C H II region shown at 20 cm (Heywood et al., 2022; Yusef-Zadeh et al., 2022a), 8 µm (Stolovy et al., 2006), 24 µm (Carey et al., 2009), and 70 µm (Molinari et al., 2010), overlaid with 214 µm magnetic field pseudovectors from Paré et al. (2024). The bright source in the top left corner all all four panels is the Sgr C H II region. The orange segments in the lower right panel are the 850 µm magnetic field detections from Lu et al. (2024). The location of an IRDC near $l=359.32$, $b=-0.17$ is labelled by the orange arrow in the top right panel.
@@
Unlike most NTFs in the CMZ, which are perpendicular to the Galactic plane, the G359.33-0.15 filaments are much closer to being parallel to the Galactic plane. Another of its notable characteristics is that it is apparently interacting with an infrared dark cloud (IRDC) at its southwestern end (in the top two panels in Figure 11 at near $l=359.32$, $b=-0.17$, and labelled by orange arrow in the top right panel), and that interaction has led to a strong morphological bunching and brightening of the radio emission at the interface (see Fig. 3k of Yusef-Zadeh et al. (2022b) for additional radio detail). The interaction has also led to the production of a slightly extended X-ray source at or very near that location (Yusef-Zadeh et al., 2007, their designation: G359.32-0.16), raising the possibility that this location is where the relativistic electrons that illuminate the G359.33-0.15 filament are produced, by analogy with G359.1-0.2 (Uchida et al., 1996) or G0.18-0.04, and the Sickle (Serabyn & Morris, 1994). This interesting structure clearly warrants further investigation. Unfortunately, no magnetic field detections have been measured along the G359.33-0.15 NTF and the IRDC with which it is interacting due to its low far-IR luminosity (Paré et al., 2024).
@@
Figure 12: The FIREPLACE 214 µm magnetic field pseudovectors around Source C, superimposed on emission maps at 20 cm (Heywood et al., 2022), 8 µm (Stolovy et al., 2006), 24 µm (Carey et al., 2009), and 250 µm (Molinari et al., 2010). The yellow segments are the 214 µm magnetic field pseudovectors reported by Paré et al. (2024). No 850 µm magnetic field detections from Lu et al. (2024) were reported near Source C. The orange arrow labels the location of the radio filament near source C.
@@
Source C contains a short radio filament with a bright southern end that ends abruptly at $l=359.390$, $b=-0.077$ in the MeerKAT/L-Band image (labelled in the top left panel of Figure 12). To the northeast of that point, there are two adjacent and parallel filaments, the upper one of which appears to undergo a sudden northward deflection at $l=359.405$, $b=-0.06$, above which its brightness increases. The deflection point of this filament is located within a far-IR emission cavity centred at $l=359.405$, $b=-0.060$, where we find no significant magnetic field detection, but north of the cavity, the magnetic field orientation is approximately aligned with the filament. If the filament is a magnetic NTF, its “deflection” of the top filament could occur where it transitions going northward from the far-IR cavity into the local molecular cloud, thereby reflecting the differing field orientations in the two media.
@@
Figure 13: Top left: The full extent of the Sgr C NTF in radio continuum (Heywood et al., 2022). The orange contours are the FIREPLACE 214 µm emission at (3, 5.25, 7.5, 9.75, 12) Jy/pixel from Paré et al. (2024). Note the faint southern extension of the bright NTF bundle to the north. Top right: The faint southern “tail” of the Sgr C NTF, and source G359.467-0.17, displayed at wavelengths of 20 cm (Heywood et al., 2022) and 8 µm (Stolovy et al., 2006). Bottom: The extended 20 cm radio emission towards the south of the Sgr C H II region, from Heywood et al. (2022), 8 µm (Stolovy et al., 2006), 24 µm (Carey et al., 2009), and 70 µm (Molinari et al., 2010), overlaid with the 214 µm magnetic field pseudovectors from Paré et al. (2024).
@@ 5.2 The Sgr C H II Region
To shed more light on the properties of the Sgr C H II region, we use the [C II] line analysis conducted by Riquelme-Vásquez et al. (2025, submitted) using the upgraded German Receiver for Astronomy at Terahertz Frequencies (upGREAT) instrument aboard the SOFIA telescope. The [C II] line traces the warm ionized and partially ionized media exposed to UV photons from the HII region. Figure 14 displays the C II velocity channel image at $-41$ km s-1 (Riquelme-Vásquez et al., 2025, submitted) overlaid by the FIREPLACE 214 µm magnetic field pseudovectors for comparison with the morphology of the C II shell.
Figure 14: The CMZ in [C II] Survey image of the $-41$ km s-1 velocity channel observed using SOFIA/upGREAT (Riquelme-Vásquez et al., 2025, submitted). White segments are the 214 µm pseudovectors made from data reported by Paré et al. (2024). Pink stars indicate the positions of two Wolf-Rayet stars reported by Geballe et al. (2019). The cyan schematics mark the approximate location and extent of the Sgr C H II region and the Sgr C NTF.
This particular velocity channel is chosen to best highlight the cavity structure revealed by the [C II] line. In the immediate vicinity of the Sgr C H II region, the 850 µm magnetic field measurements reported by Lu et al. (2024) are largely consistent with the 214 µm ones shown here.
Figure 14 reveals what appears to be a double-cavity morphology associated with the Sgr C H II region, with the western cavity being larger in size. The [C II] shell surrounds and overlaps with the outer perimeter of the H II region appearing in radio emission (see Panel #9 in Figure 7), presumably identifiable as the ionization front. The 214 µm magnetic field orientation is mostly tangential to the cavities, except in the southeast direction because of the dense foreground head-tail cloud. The [C II] emission in this shell is found to deviate somewhat from circular symmetry, with the strongest emission in the northwest and near the G359.44-0.10 EGO. Together with the tangential magnetic field morphology, these observations resonate with the conclusion found in the simulations of Arthur et al. (2011), that the expansion of an H II region creates a tangential magnetic field orientation around its border. Moreover, the [C II] cavities are separated by a “bridge” of faint [C II] emission centred at $l=359.455$, $b=-0.090$. Only two significant 214 µm pseudovectors are present toward the “bridge,” and both are parallel to it.